### v1

# Future-as-Label: Scalable Supervision from Real-World Outcomes

Abstract
Many real-world prediction problems lack labels observable at prediction time, creating a temporal gap between prediction and outcome that yields supervision only after events resolve. To address this setting, we extend reinforcement learning with verifiable rewards to temporally resolved real-world prediction, and use it to train language models to make probabilistic forecasts under causally masked information with retrospective evaluation using proper scoring rules. Supervision is derived solely from post-resolution outcomes, preserving delayed-reward semantics. On real-world forecasting benchmarks, Qwen3-32B trained using Foresight Learning improves Brier score by 27% and halves calibration error relative to its pretrained baseline, and outperforms Qwen3-235B on both constructed future-event prediction tasks and the Metaculus benchmark despite a 7× parameter disadvantage.
Machine Learning, ICML
1 Introduction
Reinforcement learning with verifiable rewards has recently emerged as an effective approach for improving language models in domains such as mathematics, code generation, and formal reasoning, where correctness can be checked automatically. By replacing human annotation with deterministic reward functions, these methods scale efficiently and yield strong empirical gains. However, their applicability depends on the availability of immediate, closed-form verification, restricting them to tasks where correctness can be resolved at training time. As a result, despite their success, existing RLVR approaches remain confined to a narrow class of problems defined by readily available, task-specific reward signals.
In contrast, many real-world processes evolve over time and resolve to objective outcomes that are independent of the model. These outcomes, such as the conclusion of an election or the decision in a court case, are publicly observable and verifiable after the fact. This temporal structure induces a natural asymmetry between the information available at prediction time and the information revealed at resolution, creating a setting in which predictions can be evaluated retrospectively without relying on contemporaneous labels.
Our goal is to translate this temporal structure into a scalable learning framework for language models. Building on prior work on Foresight Learning (Turtel et al., 2025), we formalize learning from real-world temporal streams by grounding supervision in event resolution. The key constraint is causal: at prediction time t, the model is restricted to information available up to t, while evaluation is deferred until the corresponding outcome is realized. This formulation extends reinforcement learning with verifiable rewards beyond closed-world tasks with immediate feedback to settings where correctness is determined only after external, real-world resolution.
We adopt a reward-based objective that frames prediction as a stochastic decision evaluated retrospectively after outcome resolution. In contrast to supervised fine-tuning, which fits fixed targets, Foresight Learning optimizes over sampled reasoning trajectories using only outcome-based rewards, without intermediate annotations or task-specific labels. This perspective emphasizes calibration and decision quality rather than target matching. While we focus on binary outcomes for clarity, the formulation generalizes naturally to richer outcome spaces, such as continuous, multi-class, and free-text outcomes.
In this work, we formalize learning from temporally resolved real-world events as an extension of reinforcement learning with verifiable rewards, introduce an annotation-free algorithm for learning from delayed, outcome-based supervision, and show that this approach yields substantial improvements in calibration and predictive accuracy over strong pretrained baselines.
2 Related Work
2.1 Reinforcement learning with verifiable rewards
Reinforcement learning with verifiable rewards (RLVR) has demonstrated strong results in domains with immediate, algorithmically checkable feedback, such as mathematics and programming (Wen et al., 2025) (Su et al., 2025). These settings typically involve short horizons and tightly scoped environments, which simplify credit assignment and reward attribution. Foresight Learning extends this paradigm to settings where outcomes resolve only after substantial temporal delay and outside the model’s control.
2.2 LLM-based forecasting and static supervision
Recent work applies large language models to forecasting real-world events using prompting, retrieval, ensembling, and supervised fine-tuning over historical questions (Halawi et al., 2024). In particular, Halawi et al. generate multiple candidate reasoning–prediction pairs and then use realized outcomes (via Brier score) to select high-performing outputs for supervised fine-tuning. While outcome information is therefore used for offline data curation, this learning remains non-interactive. Foresight Learning differs by incorporating outcome resolution directly into the training loop as reinforcement signals.
2.3 Model-based judges and endogenous rewards
Another line of work uses language models as evaluators or judges to provide scalable feedback in settings where objective reward functions are unavailable (Liu et al., 2025b). Such approaches enable efficient supervision, but the resulting rewards are model-mediated and derived from re-evaluating the same information available to the predictor, which can propagate the biases and limitations of the judge model.
Foresight Learning differs in the source of supervision rather than the evaluation mechanism itself. While language models may assist in outcome resolution (e.g., assessing free-text evidence), the resolver has access to information that is causally unavailable at prediction time. Rewards are therefore grounded in externally resolved outcomes rather than alternative interpretations of the same input, ensuring that supervision reflects genuinely new evidence revealed over time.
3 Method
We consider settings where supervision is provided by the eventual resolution of events rather than contemporaneous labels. At prediction time t, the model observes only information available up to that cutoff and predicts whether an event will occur by a later time s >t. Although training is performed on events whose outcomes are already known, inputs are causally filtered to exclude post-t information, and rewards are computed solely from outcome resolution at time s, preserving the temporal asymmetry of prediction by construction.
3.1 Learning formulation
Each episode corresponds to a single future-event prediction.
Predictor and resolver roles.
Foresight Learning decomposes learning from temporal streams into two roles with asymmetric information access:
  • The predictor is the language model being trained. At time t, it observes a temporally masked information state and produces a probabilistic prediction about a future event.
  • The resolver is an external, fixed process that determines the realized outcome once the event resolves at time s > t. The resolver is implemented using a pretrained, frozen language model that is not trained, updated or influenced by the learning process. The resolver has access to post-t information unavailable to the predictor and is used solely to resolve outcomes, not score or rank predictions.
The predictor and resolver are strictly separated: the predictor never observes resolution information, and the resolver does not observe model outputs or training dynamics. Learning is driven by the information gap between the predictor’s masked view at time t and the resolver’s unmasked view at time s.
State.
The state consists of all information available up to time t, including relevant dated text and a natural-language specification of an event guaranteed to resolve by time s > t. The predictor operates under a masked information state, with all post-t information causally excluded by construction.
Action.
Conditioned on the state, the policy samples an internal reasoning trajectory terminates in a probabilistic prediction $p\in(0,1)$, represented as a scalar value rather than a generated token. Only this numeric probability is exposed to the environment.. Formally, the action is the emitted probability; the trajectory is an internal stochastic computation optimized via policy gradients.
Reward.
Once the event resolves, a terminal reward is assigned using the log score:
$\text{Reward}=y\cdot\log(p)+(1-y)\cdot\log(1-p)$
where $y\in\{0,1\}$ is the realized outcome. This strictly proper scoring rule incentivizes calibrated probabilistic predictions and provides a continuous learning signal under uncertainty.
Outcome determination is performed by a separate resolver that observes the unmasked future. The resolver has access to post-t sources unavailable to the predictor and is used solely to verify whether the event occurred. Each episode terminates after outcome resolution; there are no intermediate rewards.
Although the terminal reward takes the form of a proper scoring rule, this learning setup is not simply supervised likelihood training. In expectation, optimizing this reward corresponds to maximizing the log-likelihood of realized outcomes conditioned on the information available at prediction time. However, the learning problem is structured differently: the predictor acts under a causally masked information state without access to outcomes, and training optimizes a stochastic policy over reasoning trajectories whose quality is evaluated only after outcome resolution. Credit assignment is performed via policy gradients on sampled trajectories rather than by directly differentiating a likelihood objective, preserving the decision-theoretic structure of acting under asymmetric information.
This formulation treats prediction as a stochastic decision evaluated retrospectively after outcome resolution, distinguishing it from supervised likelihood training even though the reward takes the form of a proper scoring rule.
3.2 Objective and optimization
The objective is to maximize expected terminal reward under outcome-based supervision. In this regime, single-sample policy gradients exhibit high variance due to sparse terminal feedback and intrinsic uncertainty in event outcomes. To address this, we optimize using Group Relative Policy Optimization (GRPO), as formulated by (Liu et al., 2025a).
For each state, the policy samples a group of K trajectories, each producing a probabilistic prediction. After outcome resolution, a reward is computed for each trajectory. We define a group-relative advantage by subtracting the mean reward within the group:
$\text{Advantage}(\tau_{i})=\text{Reward}(\tau_{i})-\frac{1}{K}\sum_{j}\text{Reward}(\tau_{j})$
Policy updates maximize the expected advantage-weighted log probability of each trajectory under the current policy.
By comparing trajectories generated under identical pre-t information, GRPO reduces variance from outcome noise and stabilizes learning when supervision is provided only through terminal outcomes. Gradients are applied to all tokens in each trajectory, enabling credit assignment across extended reasoning processes even though feedback is available only at the final step.
3.3 Training protocol
We explicitly enforce a causal information constraint by applying a temporal information mask to the input stream. For each prediction time t, the predictor is restricted to observing only information timestamped at or before t, even though training is performed offline. All training events resolve strictly after the pretrained model’s knowledge cutoff, ensuring that realized outcomes cannot be encoded in the model’s parametric memory. All post-t information - including sources required to determine the outcome - is withheld during prediction and policy optimization. Outcome verification is performed by a separate resolver with access to the unmasked stream. This preserves a strict causal separation between observation, action, and verification throughout training.
Figure 1: Overview of Foresight Learning
4 Experimental Setup
4.1 Dataset construction
We construct a future-event prediction dataset designed to preserve a strict temporal separation between prediction and verification. The pre-cutoff information state consists of an English-language news corpus aggregated from publicly accessible outlets (e.g., international newspapers, wire services, and financial news sites). Articles are timestamped using publisher-provided publication times, normalized to UTC.
For each example, we freeze the news corpus at a cutoff time t and generate a binary question about an event expected to resolve strictly after that cutoff, using only information available prior to t. The cutoff is defined with respect to publisher-reported publication timestamps; articles with missing or ambiguous timestamps are excluded to prevent temporal leakage. Generated events span multiple domains, including politics, economics, and corporate actions.
To prevent information leakage, model inputs are constructed exclusively from sources published at or before time t. Outcome verification uses independent post-t sources that are not included in the model’s input context. Event outcomes are resolved automatically by a separate, frozen large language model (Gemini-2.5-Flash) with access to a broader pool of post-cutoff news and archival sources, and used solely to determine whether an event occurred. The resolver does not observe model outputs or training dynamics; as a result, resolution errors introduce noise but do not induce endogenous reward signals. Examples that cannot be resolved with high confidence are discarded. Each event is assigned a resolution time s, defined as the earliest dated source supporting the resolved outcome.
All questions and outcomes are generated prior to training, enabling fully offline optimization while preserving the temporal and causal structure of real-world prediction.
4.1.1 Dataset Statistics and Splits
The full dataset contains 5,620 binary prediction examples. Of these, 5,120 examples are used for training, and 500 examples are held out as a temporally disjoint test set constructed using the same event-generation procedure. In addition, we evaluate on a second, independent test set consisting of 293 human-written forecasting questions from Metaculus, which are never used during training or data construction.
We intentionally do not construct a validation set: models are trained on all available pre-test data using a fixed training procedure, without early stopping or model selection based on held-out examples. Training data consists of predictions made as of July 1, 2024 through January 30, 2025. Both test sets consist exclusively of predictions made on or after February 1, 2025, ensuring strict temporal separation between training and evaluation. The dataset described in this section is released publicly to support reproducibility and future research.
4.1.2 Task characteristics
Prediction horizons range from days to several weeks, allowing learning across varying outcome horizons. Although outcomes are discrete, supervision and evaluation are based on continuous probabilistic scores, enabling analysis of accuracy and calibration under increasing temporal uncertainty.
4.2 Models and training
We fine-tune a Qwen3-32B language model with explicit reasoning enabled. Conditioned on an information state, the model generates a reasoning trajectory that terminates in a probabilistic prediction expressed explicitly at the end of the output. Parsed probabilities are constrained to the interval [0.001, 0.999] for numerical stability.
Training is performed using GRPO. For each event, the model samples four independent trajectories, each producing a probabilistic prediction. After outcome resolution, a log-score reward is computed for each trajectory, and relative advantages are obtained by subtracting the per-group mean reward. Policy updates increase the relative likelihood of higher-reward trajectories. Training uses batches of 32 events, with prediction horizons mixed within each batch.
4.2.1 Baselines
We compare Foresight Learning to baselines that operate under identical temporal constraints and produce probabilistic predictions in the same output format, isolating the effect of learning.
Prompted forecasting: the base Qwen3-32B and Qwen-235B models are prompted to produce probabilistic predictions without task-specific fine-tuning. This baseline measures forecasting performance without learning from outcome resolution.
Ensembling: multiple independent predictions are generated per event and averaged to assess gains from sampling and aggregation without parameter updates. This control tests whether improvements can be explained by variance reduction alone.
4.2.2 Evaluation metrics
We evaluate models based on the quality of probabilistic predictions. We report the log score used for training, the Brier score, which measures squared error between predicted probabilities and outcomes, and calibration, assessed via expected calibration error (ECE) over 10 discretized probability bins measuring empirical outcome frequencies as a function of predicted confidence.
5 Results
We evaluate Foresight Learning on two held-out test sets: (i) a synthetic future-event benchmark of 500 questions constructed under strict temporal controls, and (ii) an external benchmark consisting of 293 binary forecasting questions from Metaculus. Performance is evaluated using proper scoring rules and calibration metrics.
Table 1 compares four inference regimes: (i) Qwen3-32B prompted for a single forecast, (ii) Qwen3-32B prompted for seven independent forecasts with the median taken as the final prediction, (iii) Qwen3-235B prompted for a single forecast, and (iv) the Foresight-trained model prompted once. Repeated prompting and median aggregation provide modest improvements over single-sample prompting but do not match the gains from training on resolved outcomes. Notably, the Foresight-trained 32B model outperforms both the ensemble-style baseline and the substantially larger 235B model across all metrics, indicating that the improvements stem from the training objective rather than increased sampling or model scale.
Performance gains persist on the Metaculus benchmark, which consists of independently authored questions outside the synthetic benchmark distribution. One possible contributing factor is that Metaculus questions often concern higher-salience events with broader public coverage, providing richer information at prediction time. While this hypothesis requires further study, the results indicate that learning from externally resolved outcomes generalizes beyond the specific data construction process used for training.
Taken together, these results support the central premise of Foresight Learning: incorporating outcome resolution directly into the training objective yields more accurate and better-calibrated probabilistic forecasts than prompting or sampling-based baselines alone, even when compared to substantially larger pretrained models.
Table 1: Forecasting performance on synthetic and real-world benchmarks
Model Log $\uparrow$ Brier $\downarrow$ ECE $\downarrow$
Metaculus
Qwen3-32B -0.7210 0.2472 0.2175
Qwen3-32B Ensemble -0.7000 0.2390 0.2289
Qwen3-32B-RL (160) -0.5738 0.1793 0.1042
Qwen3-235B -0.6828 0.2111 0.1905
Synthetic future-events
Qwen3-32B -0.7166 0.2432 0.1732
Qwen3-32B Ensemble -0.7045 0.2481 0.1864
Qwen3-32B-RL (160) -0.5978 0.1979 0.0598
Qwen3-235B -0.7138 0.2260 0.1695
Figure 2: Model calibration and accuracy metrics versus training steps on Metaculus (top) and synthetic future-events (bottom). Shaded regions show 95% bootstrap confidence intervals. Metrics are log score (↑), Brier score (↓), and expected calibration error (ECE; ↓). Performance improves monotonically with training.
Figure 3: Brier scores (↓) for different models on Metaculus and synthetic future-events benchmarks. Foresight Learning consistently outperforms both the base and larger pretrained baselines.
6 Discussion and Conclusion
This work studies a supervision regime in which feedback is provided by the eventual resolution of real-world events rather than contemporaneous labels or proxy objectives. By optimizing probabilistic predictions retrospectively using proper scoring rules, Foresight Learning aligns training with the temporal and causal structure of forecasting under uncertainty.
Empirically, learning from outcome resolution improves probabilistic forecasting performance relative to a strong pretrained baseline, with consistent gains in accuracy and calibration on both synthetic future-event datasets and the independently authored Metaculus benchmark. Notably, Foresight Learning materially outperforms a substantially larger same-generation model on real-world forecasting tasks.
A key benefit of outcome-based supervision is improved calibration. Because rewards are assigned only after outcomes resolve, overconfident incorrect predictions incur large penalties, while appropriately uncertain predictions are penalized less severely. This learning signal encourages inference strategies that balance evidence aggregation with uncertainty estimation, whereas sampling-based heuristics such as ensembling reduce variance without modifying the underlying prediction policy.
Relative to prior reinforcement learning with verifiable rewards, Foresight Learning operates in open-world domains with sparse and delayed feedback. Trajectory-level, group-relative optimization enables stable credit assignment under long and variable horizons by comparing alternative predictions generated under identical informational constraints and evaluating them retrospectively after outcomes resolve.
This work has several limitations. Training is performed offline on resolved events; while deployment-time feedback loops are under active exploration, they are not evaluated in this study. Event specification and outcome resolution rely on automated pipelines that may introduce biases or coverage gaps, and the current experiments focus on binary outcomes. Extending the framework to richer outcome spaces and fully online settings remains an important direction for future work.
Overall, Foresight Learning demonstrates that effective supervision can arise directly from chronologically evolving real-world data. By incorporating outcome resolution into the training objective, the framework points toward a broader role for outcome-based supervision in extending verifiable reward-driven learning beyond closed-world tasks and toward open-ended, real-world decision-making.